OpenStudy (anonymous):

Find the area of a kite with diagonals measuring 6 cm and 18 cm?

OpenStudy (ganpat):

Area of Kite is given by = (product of diagonal )/ 2 A = (6 * 18) / 2 = 54.. sq cm..

OpenStudy (anonymous):

where did you get 54 from? sorry, i really don't remember any of this

OpenStudy (anonymous):

so 18 x 6 = 108, and i divide that by 2?